56) Minneapolis business Rogue Chocolatier sells specialty chocolate bars with a high cocoa content. Rogue can produce 15 chocolate bars per day with one employee, 29 with 2, 35 with 3 and 40 with 4 employees. Which statement is TRUE?

A) Rogue's marginal product curve decreases initially and then increases.

B) Rogue's marginal product curve initially increases and then decreases.

C) Rogue's total product curve increases at an increasing rate.

D) Rogue's marginal product curve decreases continually.

57) In 2008, Agriculture Secretary Ed Schafer announced that Chile's Livestock and Agricultural Service approved the U.S. inspection, control and certification systems for poultry, allowing these products to enter the Chilean market. The effect of this change in Chilean policy in the Chilean poultry market is to

A) decrease Chilean total surplus.

B) increase Chilean consumer surplus.

C) increase Chilean producer surplus.

D) have no effect in the Chilean poultry market.

58) In the long run, perfectly competitive firms make zero economic profit. This result is due mainly to the point that a perfectly competitive market has

A) firms with perfectly elastic market demand.

B) price taking by the firms.

C) no barriers to entry and exit.

D) few buyers and sellers.

59) Dell uses outsourcing, that is, Dell buys the components of the computers it produces from other firms. This is an example of

A) a partnership.

B) market coordination.

C) a command system.

D) firm coordination.
Answer #1

56) D) Rogue's marginal product curve decreases continually.
(First worker adds 15 units; 2nd worker adds = 29 - 15 = 14 units; 3rd worker adds = 35 - 29 = 6 units; 4th worker adds 40 - 35 = 5 units. So, MP is decreasing always.)

57) B) increase Chilean consumer surplus.
(This will cause an increase in supply, which reduces price and decreases quantity so consumer surplus increases.)

58) C) no barriers to entry and exit.
(Due to no barriers, firm will enter and exit until each existing firm earns zero economic profits.)

59) B) market coordination.
(This is an example of market coordination.)